Reports 1-1 of 1 Clear search Modify search
MIF (ITF Control)
masayuki.nakano - 5:46 Sunday 26 January 2020 (12639) Print this report
Out first recycling interferometer!!

Finally, we succeeded in locking the power-recycled Fabry-Perot Michelson Interferometer (PRFPMI)!!

The PRFPMI has been controlled with IR signals and the lock acquisition is almost automated. PRFPMI is locked within 5 min after the lock-loss and the lock duration is more than 10 min (I kicked for the measurement, and did not try to lock for long).
We have not calibrated the signal yet. We gonna calibrate it as soon as possible.
The detail is as follows:
  • The signal used for each DoF is:
    PRCL: POP45I
    DARM: AS17Q
  • Lock acquisition procedure is as follows:
  1. Lock ALS CARM and DARM
  2. Bring it to the resonance by sweeping the summing node offset and the output bias of the PDH-Y servo.
  3. Lock PRMI by using 3f (REFL51Q for MICH and REFL135I for PRCL)
  4. Lock DARM by mixing TRNORM AS17Q to ALS DARM signal
  5. Lock CARM by mixing TRNORM REFL45I to analog ALS CARM signal without offset.
  6. Remove ALS signals.
  7. Handoff the vertex control to the 1f signal shown above.
  • The TRANS power of each cavity is 120 and 100 for Xarm and Yarm, respectively. These numbers are normalized by the TRANS power with the single-arm and PRM misaligned. If the PRM transmissivity is assumed as 10 %, the PRG is 10~12. We are not sure what causes this asymmetric TRANS power.
  • Guardian is ready for the PRFPMI lock, except for the transition from MICH 3f to 1f. This transition is not really stable and I succeeded only several times. Also, the stability of the lock with the 1f signal is not good and lost lock in ~10sec. We need an investigation. The input matrix component for MICH1f is -0.0015.
  • The SNR of POP45I and REFL135I seemed not really different. This might be because the laser power on the PD is low as ~1 mW. We need to check if there is room to increase it.
  • I only measured the DARM, the PRCL OLG, and no optimization  I will measure several OLGs and tune the gain and filter shape.
Build-up signals(Top, IR X TRANS (left), Y TRANS(mid), build-up sideband of f2(right). DC PD signals(mid), Feedback signals (bottom).

Images attached to this report
Comments to this report:
osamu.miyakawa - 7:53 Sunday 26 January 2020 (12640) Print this report
Very very good news! I wanna see the sensitivity sensitivity ASAP.
Search Help